# Bridging the Gulf of Envisioning: Cognitive Design Challenges in LLM Interfaces

### v1

@@ 2.3. Role of Design in Mapping Intentions to Actions
Until now, HCI approaches have successfully interwoven intentions and actions during interaction design. Through human-centered design approaches, interaction designers strive to ensure that every potential action available on an interface aligns intuitively with the users’ goal-based intentions (Wood, 1997; Norman, 2014). Even for complex cognitive goals such as reading, writing, reasoning, and sense-making, researchers have developed interactive interfaces by understanding the underlying cognitive process model (e.g., (Kuznetsov et al., 2022; Subramonyam et al., 2020; Endert et al., 2015)). For instance, in texSketch (Subramonyam et al., 2020), they build interactions around the cognitive processes in reading, such as selection, organization, and integrated comprehension. This alignment fosters a symbiotic relationship where users feel the system is an extension of their thought processes, making the interaction feel natural and efficient.
@@
Second, design also plays a critical role in helping users formulate a mental model of the system, and through it, learn to anticipate system outcomes. In conventional HCI, designers invent a conceptual model – an abstract representation or framework to communicate how a system operates – based on human-centered design practices (Norman, 2014). For instance, the designer may analyze users’ past experiences, and prior knowledge, and draw from familiar analogies and metaphors (e.g. the desktop metaphor) to shape the presentation and behavior of the system (Neale and Carroll, 1997). According to Hutchins (Hutchins, 1987), conceptual models function as metaphors on several levels. At the broadest level, they’re shaped by the users’ primary goals. In a word-processing tool, this could mean a “blank page” metaphor aligning with users’ primary goal of creating a document. Moving in more granularly, metaphors at the interaction level inform users about their computer interactions and remain consistent across tasks (e.g., a “clipboard” for copying and pasting). Lastly, metaphors at the task domain level provide an understanding of how tasks are organized and structured, such as the use of tabs or headers and footers to indicate document structuring during writing. By establishing this foundation during design, users can more easily form a coherent internal representation of the system’s operations and functionalities. This representation enhances their ability to plan and execute complex tasks while seamlessly aligning their goal intentions with the appropriate actions.
With LLM systems, the link between user intentions and system actions is less clear, and users lack a conceptual understanding of how an LLM functions. Given the very different nature of transformer processing in machine learning, even experts are unable to predict how inputs determine outputs in LLM systems. Consequently, LLM interactions are challenging to design based on the conventional interaction approaches in HCI.
2.4. LLM Interactions as Intentions Without Actions or System Mental Models
LLMs represent a significant leap in the evolution of natural language processing (NLP) capabilities. On the surface, they seem to understand natural language inputs, engage in conversational interactions, and can carry out a wide range of generative tasks, including writing, coding, and explaining. From the perspective of interactive system design, LLMs obviate the need for structured interfaces with set actions for implementing intentions in favor of unconstrained use of natural language (note that we address specific interface designs for LLMs in Section 4). However, the paradox is that language use as the interface makes LLM interactions more challenging for users. In LLMs, the formulation of intentions lacks a provided mapping onto explicit interface actions. This makes it harder to interact because users have to do more cognitive work to envision a process model for the task, identify system instructions, and evaluate the desirability of outcomes.
More importantly, users do not have good system mental models mapping action specifications to the system-produced outcomes. LLMs are trained on massive corpora of data and optimize the probability of words based on appearances in that corpus. As a mathematical solution based on word occurrence, we lack a comprehensive understanding of why they work. The machine learning community has postulated high-level ideas of how LLMs are able to accomplish different natural language tasks; for example, from a pre-training perspective, it has been hypothesized that LLMs can generalize well to new tasks due to variation in tasks during training (Radford et al., 2019). Specific prompting techniques such as zero-shot prompting, few-shot prompting, and chain-of-thought prompting are emergent capabilities that are also attributed to the model’s scale without providing an understanding of why they work (Radford et al., 2019; Brown et al., 2020; Wei et al., 2022). Without system mental models, users cannot effectively predict outcomes or explain interactions with LLM models. This often leads users to pose ambiguous and ill-defined queries and yet expect the LLM to understand and respond appropriately.
To address this paradox, in the next section, we propose a new HCI approach for the design of LLM interfaces to support a process we term envisioning. This proposed envisioning process directly addresses the transformation in systems from structured task interfaces to the natural language interfaces of LLMs.
3. The Revised Model of Interaction for LLM
Figure 2. In the context of Norman’s seven-stage model action, we highlight what is missing during human-LLM interactions. Further, there are three pathways to interactions: (1) directly state their goal to the LLM, (2) formulate their intentions and provide them to the model through prompt engineering, and (3) taking the LLM output and transitioning to a dedicated interface and system (e.g., switching from ChatGPT to a Word Processor based on an LLM generated draft).
When a system mental model of output production is lacking, how should a user interact with an LLM to meet a goal? One interpretation of this question is that the user is not interacting with the LLM in the conventional sense; instead, they are “acting as if” the LLM can directly understand their intentions, an instance of the pervasive phenomenon of treating computers and their computational processes as if they were people (Reeves and Nass, 1996). We will later discuss the conversational nature of this interchange and how it differs from human-human conversations. We begin here by focusing on a single input to the LLM, which involves the articulation of intentions, by imagining the output, to formulate the prompt for the LLM. We call this process envisioning. When compared to the mental model simulation of system functions we discussed in Section 2, envisioning requires end-users to create and operate their own model of cognitive task processes to accomplish a goal.
3.1. Cognitive Task Processes
A generalized view of cognitive task processes for ‘generative’ tasks, such as writing, creating new ideas, coding, and reasoning, is an iterative process of planning intentions, executing, and then evaluating, with new intentions generated with each iteration (Bamberger and Schön, 1983). In other words, generative tasks typically take the form of repetitive cycles of “plan a little” and “do a little.” In cognitive task processes, planning involves identifying and organizing sub-goals, intentions, and actions (Miller et al., 2017; Pea, 1982). The process creates a path through a space defined by the current goal state, a desired goal state, and available actions to bridge the gap (Newell et al., 1972). For instance, a means-ends analysis process discerns disparities between a current and a desired state, and selects actions to decrease their distance. This includes searching the action space, back-tracking from goal to current states, mental simulation to gauge action viability, and branching and chunking to make the planning process more efficient given human memory constraints (Miller, 1956; Decety and Grèzes, 2006). Next, an execution phase turns intentions into actions. In addition to “doing” the task, users also coordinate other activities, such as monitoring for errors and making real-time adjustments based on feedback. During an evaluation phase, users are engaged in a comparative judgment of the actions’ outcome in relation to a goal state. Evaluation also includes self-regulation and monitoring, metacognitive awareness, and attribution, which is determining causes of outcomes in order to plan subsequent intentions and actions (Brown, 2017; Schraw and Dennison, 1994; Wittrock, 1989).
@@ 3.1.1. A Cognitive Task Process for Writing Tasks
A cognitive process model for writing posits writers first define the task environment, including the audience, the purpose, and tools (Hayes, 2013). Next, the act of writing itself involves three intertwined cognitive stages, including planning, translating, and reviewing. The planning phase provides a mental roadmap for writing. It involves retrieving ideas or content related to the writing topic, articulating objectives such as the intent to inform, persuade, or entertain, and structuring the ideas in a coherent manner by deciding which points to introduce first, and how to transition from one idea to the next. Translating is the act of navigating the planned roadmap and putting those ideas into words by converting cognitive representations into linguistic expressions. Third, after a segment of text (or the entire piece) is written, writers review it to determine whether it aligns with their initial goals, such as whether the content is coherent, compelling, and clear. Based on this evaluation, they may choose to revise the text by restructuring entire paragraphs, adding new information, or deleting redundant or suboptimal content. Crucially, this cognitive process during a writing task is not simply a linear progression from one stage to the next, but an iterative, dynamic process of constant refinement and reflection as writers respond to ongoing assessment of their output.
3.2. Envisioning LLM Interaction and Associated Gaps
@@
Based on this understanding of cognitive task processes, let us revisit the goal of producing the Wikipedia Article on Chocolate. Our aim is to identify different pathways to accomplish the goal using LLMs and identify the associated challenges that a user might face. The most straightforward path is to simply state the goal to the LLM (i.e., with no envisioning) as shown in Figure 2-1, such as, “Write me a Wikipedia Article about Chocolate.” Such an approach aligns with the theory of computational rationality, in which interactive behaviors fundamentally depend on the principle of expected utility (Oulasvirta et al., 2022). At its core, users tend towards behaviors that maximize expected utility, bounded by specific constraints. These limitations can emerge internally from cognitive or bodily capacities, or externally from the environment. Of course, given that LLMs are trained on Wikipedia data, the resulting output may satisfice a user’s goal (Simon, 1956; Tidwell, 2010).
@@ 3.2.1. Intentionality Gap:
In this interaction with the LLM, the user states their goal directly, and forgoes any cognitive task processes involved in doing the task themselves (i.e., planning what to write, executing it, and evaluating it). Instead, they start with the evaluation of the LLM-generated text. Does it meet their goal? This evaluation is cognitively demanding because no “bridging” steps between goal and output are provided. By skipping directly from goal to evaluating output, users may lack a cohesive understanding of the task context and required content, hindering their ability to make comprehensive judgments about the adequacy of the output. Drawing from philosophical terminology, users encounter an intentionality gap while assessing texts generated by LLMs. In philosophy, intentionality refers to the capacity of mental states to be about, directed at, or represent something in the world outside of themselves (Searle, 1983). In the LLM context, users may not have a nuanced “aboutness” of the specific intentions or contexts they are considering and, therefore, may not have a clear sense of what the content should be “about” from their own perspective. Consequently, without internal cognitive benchmarks of the generation process for comparison, they may become over-reliant on the LLM and avoid evaluation against goals. Further, engaging in subsequent planning to identify and rectify any deviations from a desired outcome can be challenging. In other words, users producing only a goal statement will have difficulty answering questions such as “Can the LLM do better?” and “How can the LLM do better?”
@@
Note that this intentionality gap is different from the gulf of evaluation in Norman’s model (Norman, 1986). With traditional HCI tasks, success is defined as binary – a task was either completed successfully or not. Feedback is often immediate, as with a direct manipulation user interface in which “users can immediately see if their actions are furthering their goals (Shneiderman, 1982).” The evaluation gulf stems from ambiguous system feedback at the interface level, which may make it challenging to perceive and interpret the produced outcome. Addressing the gulf of evaluation involves improving system feedback, representation, and visibility. In contrast, the intentionality gap lies at the cognitive level due to the user’s failure in creating clear intentions. Bridging this intentionality gap with LLMs may require interfaces scaffolding clearer task processes and contextual understanding in order to evaluate the generated text.
Rather than stating the goal as the input to an LLM, a second interaction pathway (Figure 2-2) for the end-user is to envision the cognitive processes involved in generation in order to provide more specific intentions to the LLM (i.e., envisioning). This is essentially the same as the emerging proposals for prompt engineering strategies (discussed in Section 7.2). For instance, in writing the Wikipedia article, the user may engage their planning processes to develop intentions about the topics to include in the generated text, its structure, constraints about format and length, source references, tone and writing style, etc. The more developed and specific user’s specified intentions, the better the LLM’s output will satisfy the goal. (This is not because they are developing a functional model of the LLM’s, but because the LLM has been trained on texts written by humans with similar functional approaches to cognitive tasks.) However, applying this in practice is not easy. To craft a more detailed prompt for the LLM, the user has to mentally engage all their planning and evaluation task processes to generate the intermediate intentions but without actually executing them. Without execution, envisioning good intention specifications can be challenging and place heavy demands on cognition. For instance, telling the LLM what to include in a section without actually writing may be harder because the intentions are not yet realized on the document to provide the opportunity for feedback. This is akin to asking someone instructions for making a peanut butter and jelly sandwich without actually making it [^1] The EXACT INSTRUCTIONS CHALLENGE video: https://youtu.be/Ct-lOOUqmyY due to the demands of generation without the feedback from execution. More to the point, advance intention setting exemplifies the gap in capturing the user’s intentionality directly, potentially leading the LLM to produce outputs diverging from the user’s intent.
Further, the absence of system mental models and interface-supported actions introduces two more gaps in the process of envisioning: (1) capability gap, and (2) language gap. We discuss them below.
3.2.2. Capability Gap:
Recall from our earlier definition that an intention includes specifications about how to execute task processes. The capability gap concerns the users’ inability to formulate “how” to implement their intentions. Defining when, where, and how one wants to take action on intentions requires added cognitive effort (Patalano and Seifert, 1997; Seifert and Patalano, 2001), but has been shown to enhance the rate of successful execution (Gollwitzer and Oettingen, 2020). While LLMs, with their extensive training data, can theoretically understand and generate a wide variety of bespoke task content, their very strength can also be a source of complexity for the user. Take, for instance, entering a prompt for summarizing a 2-page text in 1 page. For a human writer, this task involves a myriad of choices about what to prioritize, which nuances to retain, and which details to omit. How can a user identify the right specification of intention for the LLM? Users are forced to use a trial-and-error approach with differing abstractions and variations of their intentions, and then iterate based on evaluating the LLM’s output. However, unlike deterministic systems where a given input always produces the same output, an LLM can produce countless variations from the same input, and there is, by definition, no one correct answer in generative tasks. This indeterminacy in LLM responses, coupled with a lack of transparency in how inputs lead to LLM outputs, can hinder users’ attempts to formulate prompts. The language gap is defined to provide greater specification of the root causes for the capability gap.
3.2.3. Language Gap:
The language gap refers to the user’s challenges in clearly and effectively expressing their intentions in the interface as written words, leading to a potential mismatch between the user’s intentions and what the LLM perceives and then produces. One benefit to the natural language interface for LLMs for end-users is a wide-open input space for specifying intentions. Unfortunately, the LLMs’ algorithms for learning from text corpora create a dependency on language precision. While human language use tolerates a wide range of expressions communicating a similar meaning, even slight changes in words can lead LLMs to significantly different outputs. It is possible that substantial experience with LLM use will improve a user’s understanding of the precise mapping between prompt contents, at least in a domain area. Users lacking an improvisational linguistic dexterity may not obtain satisfactory outcomes through trial and error iteration of prompts. That is, they may initiate a prompt with a particular phrasing and a general sense of an anticipated outcome, but upon seeing unexpected outputs – either omissions and commissions (or both) – they are likely to need iterative refinement of their prompts to achieve their goals. This linguistic sensitivity makes it difficult for users to script their intentions for the LLM upfront by anticipating potential system interpretations. Therefore, users are likely to need a multi-turn interaction with an LLM session to determine alternative specifications for their intentions that lead to better outcomes.
3.2.4. Conversational Interactions with LLMs
The two gaps mentioned above require trial-and-error in interactions with LLMs. At the surface level, this is akin to human conversation. However, there are clear differences that must be considered with LLM interactions. When two humans are conversing, there is almost always an implicit goal of building a shared understanding. The British philosopher (Grice, 1975) developed four cooperative maxims of conversational behavior to describe specific rational principles observed by people in pursuit of effective communication: (1) The maxim of quantity is to provide as much (but no more) information than required; (2) The maxim of quality values truth with adequate evidence; (3) The maxim of relation requires relevance and pertinence to the discussion; and (4) The maxim of manner avoids ambiguity by being as clear, brief, and orderly as possible. These maxims depend on each conversationalist tracking a theory of mind (Frith and Frith, 2005) regarding what their partner thinks and knows, while LLMs do not attempt to model the understanding of the user. LLMs thus far do not have a theory of mind about their human users, and human users do not have a theory of mind for how LLMs function.
Further, humans in conversation can recognize misunderstandings, ask for clarifications, and adjust their language use and level of abstraction or detail based on continuous feedback and conduct ‘repairs’ (Schegloff et al., 1977). By contrast, Human-LLM conversations are less dynamic and interactive because LLMs do not aim to adapt or provide feedback to improve users’ expressions of their goals and intentions. Further, humans can draw from a rich context in conversations by remembering past interactions, shared experiences, and broader situational contexts. LLMs lack the ability to remember interactions, do not change as a result of interactions (unless further learning model training takes place), and lack the broader experiential context of conversations. In human conversation, rephrasing a question or response conveys the same meaning. With LLMs, a user is using conversational repairs to rephrase a prompt, but this is interpreted as a distinct and unrelated query. While a human would be annoyed at hearing the same question asked in different ways, the LLM fails to recognize a user’s rephrasing as a strategy for directing it toward a desired outcome. Users without insight into the system model or understanding of how specific phrasings influence the outputs are essentially navigating a vast and somewhat unpredictable linguistic landscape. Without the ability to view all input-output mapping during exploration, users need to remember or document their interactions to discern emerging patterns in LLM input to output over time.

### v2

@@ 2.3. Role of Design in Mapping Intentions to Actions
Until LLMs, HCI approaches have successfully interwoven intentions and actions during interaction design. Through human-centered design approaches, interaction designers strive to ensure that every potential action available on an interface aligns intuitively with the users’ goal-based intentions (Wood, 1997; Norman, 2014). Even for complex cognitive goals such as reading, writing, reasoning, and sense-making, researchers have developed interactive interfaces by understanding the underlying cognitive process model (e.g., (Kuznetsov et al., 2022; Subramonyam et al., 2020; Endert et al., 2015)). For instance, in texSketch (Subramonyam et al., 2020), they build interactions around the cognitive processes in reading, such as selection, organization, and integrated comprehension. This alignment fosters a symbiotic relationship where users feel the system is an extension of their thought processes, making the interaction feel natural and efficient.
@@
Second, design also plays a critical role in helping users to formulate a mental model of the system, and through it, learn to anticipate system outcomes. In conventional HCI, designers invent a conceptual model – an abstract representation or framework to communicate how a system operates – based on human-centered design practices (Norman, 2014). For instance, the designer may analyze users’ past experiences and prior knowledge and draw from familiar analogies and metaphors (e.g., the desktop metaphor) to shape the presentation and behavior of the system (Neale and Carroll, 1997). According to Hutchins (Hutchins, 1987), conceptual models function as metaphors on several levels. At the broadest level, they’re shaped by the users’ primary goals. In a word-processing tool, this could mean a “blank page” metaphor aligning with users’ primary goal of creating a document. Moving in more granularly, metaphors at the interaction level inform users about their computer interactions and often remain consistent across tasks (e.g., a “clipboard” for copying and pasting). Lastly, metaphors at the task domain level provide an understanding of how tasks are organized and structured, such as the use of tabs or headers and footers to indicate document structuring during writing. By establishing this foundation by means of HCI design, users can more readily form a coherent internal representation of the system’s operations and functionalities. This representation enhances their ability to plan and execute complex tasks while seamlessly aligning their goal intentions with the appropriate actions.
With LLM systems, the link between user intentions and system actions is less clear, and end-users lack adequate mental models of LLMs. Consequently, LLM interactions become challenging for users. In the next section, we postulate how these constraints challenge forming intentions for interactions.
3. Envisioning Intentions in LLM Interactions
LLMs represent a significant leap forward in the evolution of natural language processing (NLP) capabilities. From the perspective of interactive system design, LLMs obviate the need for structured interfaces with preset actions for implementing intentions in favor of unconstrained use of natural language (note that we address specific interface designs for LLMs in Section 4). With a dynamic operational scope lacking explicit interface actions, how do users formulate their intentions and then express them as prompt inputs? One proposed solution is to treat LLMs as if they are human and engage in a conversation with them (Reeves and Nass, 1996; Dillion et al., 2023; Cui et al., 2023; Zamfirescu-Pereira et al., 2023b). Here, we examine this approach to intention formation by considering a specific generative task for LLMs – writing – in three aspects of human-to-human interactions: models of communication, roles and expertise, and theory of minds.
3.1. Cognitive Processes in Generative Tasks
What happens when humans perform generative tasks? Studies show that human performance of ‘generative’ tasks, such as writing, creating new ideas, coding, and reasoning, appears endlessly variable. At a high level, human task performance takes the form of repetitive cycles of cognitive processes where a change in process, failure, or success is not predictable. A general observation is task performance proceeds through cycles of “plan a little” and “do a little” (Miller, 1956; Decety and Grèzes, 2006). There is no definitive cognitive process for how to accomplish a generative task, though much of education is aimed at instruction on exactly these tasks. Models offer few distinctions among tasks despite distinctive aims, and a high degree of variation in cognitive processes is observed for the same task in different people and for the same person repeating a task.
The predominant model of generative tasks is Newell & Simon’s (1972) model of problem-solving  (Newell et al., 1972). They defined a problem space, including the current state, a desired goal state, and all available actions or operators. The process of problem-solving is defined by actions taken to bridge the gap between nearby states, describing a path toward a solution. For instance, a means-ends analysis process identifies differences between current and desired states and selects an action to decrease their distance  (Newell and Simon, 1961). The path from each state is viewed as traversing a solution space containing all possible outcomes. Problem space models describe well-defined problems where these elements are known, and branching can identify all possible combinations of actions. This approach is successful as a model for machine problem solving (Newell and Simon, 2007), but when applied to human problem solving, deterministic solution paths with well-defined goals, operators, and evaluation of options are rare (Seifert et al., 1994). Typically, great variability is observed in cognitive processes, including strategies like back-tracking from a goal to a current state, mental simulation of potential actions, and generation of novel actions.
As a consequence, cognitive models of generative tasks describe a more variable process with ill-defined initial states, goal states, and available actions. For example, a four-stage model of human problem solving describes cognitive processes of 1) problem identification, 2) planning, 3) implementation, and evaluation (Duncker, 1945). Defining a problem is far from a determinate process yet central to success. Planning is defined as identifying and organizing sub-goals, intentions, and actions (Miller et al., 2017; Pea, 1982) to achieve a goal. Next, an execution phase turns intentions into actions. In addition to “doing” the task, people coordinate other cognitive processes, such as monitoring for errors and making real-time adjustments based on feedback. During the evaluation phase, users are engaged in a comparative judgment of the actions’ outcome in relation to a goal state. Evaluation also includes self-regulation, metacognitive awareness, and value functions, as well as explaining errors to plan subsequent intentions and actions (Brown, 2017; Schraw and Dennison, 1994; Wittrock, 1989). These stages are so loosely defined that the 4-stage model is ubiquitous in accounts of thinking, creativity, design, and other generative tasks (Duncker, 1945; Howard et al., 2008). When applied to an instance of task performance, the cognitive processes are indeterminate, with the progression and order of stages varying, and stages are often independently repeated. Iteration is assumed for all stages in these models because the cognitive processes are also defined over subtasks as needed (Bamberger and Schön, 1983).
@@ 3.1.1. A Cognitive Task Process for Writing Tasks
A cognitive process model for writing offers a more specific account of the steps required for completing a task (Hayes, 2013). First, the writer defines the task environment, including the audience, the purpose, and tools. Next, the act of writing involves three intertwined cognitive stages for planning, translating, and reviewing. The planning phase creates a mental roadmap for writing. It involves retrieving ideas or content related to the topic, articulating objectives such as the intent to inform, persuade, or entertain, structuring ideas in a coherent order, and determining how to transition from one idea to the next. The translating stage involves navigating the roadmap and putting ideas into words by converting cognitive representations into linguistic expressions. After a segment (or an entire piece) is written, a third stage involves Reviewing it to determine whether it aligns with initial goals, such as coherency, compellingness, and clarity. Based on this evaluation, they may choose to revise the text through restructuring, adding new information, or deleting content. Crucially, this cognitive process during a writing task is not a linear progression from one stage to the next, but an iterative, dynamic process of constant repetition, refinement, and reflection as writers respond to their ongoing assessment of their created output. In this more specific model of cognitive processing in writing texts, the same general steps (planning, implementing, and evaluating) and prominent iteration are evident. In addition, a role for continual monitoring of what is produced and how it meets the goals suggests constant oversight of the generation process.
3.2. Generative Task Interactions with an LLM Agent
Figure 3. In the context of Norman’s seven-stage model action, we highlight what is missing during human-LLM interactions. Further, there are three pathways to interactions: (1) directly state their goal to the LLM, (2) formulate their intentions and provide them to the model through prompt engineering, and (3) take the LLM output and transition to a dedicated interface and system (e.g., switching from ChatGPT to a Word Processor based on an LLM generated draft).
@@ 3.2. Generative Task Interactions with an LLM Agent
Based on this understanding of cognitive task processes, let us revisit the goal of writing a toast for Taylor’s retirement. If recruiting a human to help, one might choose an expert who knows Taylor well, and has been to many such events, or an assistant with strong writing skills based on prior knowledge of the expertise required for the task and a theory of specific minds reflecting their knowledge states. Based on this understanding, we may directly ask the expert to “Write a retirement toast.” With an LLM agent, we can assume training on everything posted on the internet, a generation process by next-word prediction, and conversational interaction to iterate on output. Based on this assumption, the most straightforward path is to simply state the goal to the LLM in their own words, as shown in Figure 3- PATH 1. Such an approach aligns with the theory of computational rationality, in which interactive behaviors fundamentally depend on the principle of expected utility (Oulasvirta et al., 2022). That is, users tend towards behaviors that maximize expected utility within specific constraints, called bounded rationality (Simon, 1997). Constraints can emerge internally from cognitive or bodily capacities, or externally from the environment. A key element in this determination is the perception of effort required to improve the input quality and the potential gain in output quality. Given that LLMs are likely trained on databases containing toasts, the resulting output may satisfice  (Simon, 1956; Tidwell, 2010), and generic or adequate answers may be readily available.
Experienced users may learn that instead of directly stating the goal, a greater effort to formulate intentions within prompts can increase the utility by arriving at better or quicker solutions. For example, while it is possible to iterate based on output evaluation, the time to read and reformulate prompts has costs. If the iterative steps can be combined in fewer steps, taking more care in formulating intentions may be cost-effective. For instance, in creating a prompt for the LLM to write a toast for Taylor, the user may engage in planning how they would do the task by identifying topics to include, a structure and organization, constraints for format and length, tone and writing style, desirable qualities, etc. The more developed and specific the intentions, the better the LLM’s output should satisfy the user. We call this process envisioning (Figure 3- PATH 2). However, envisioning more developed intention specifications is effortful and cognitively demanding. To craft a more specific prompt for the LLM, the user must further develop their intention by mentally exploring potential plans and values as if – but not actually – executing them, and then adding what is discovered to the prompt. This often leads to taking the Path 1 approach without envisioning. If “an answer” or even a “satisfactory answer” is needed, envisioning may not be. But, if aiming for a high-quality answer, a user may better exploit the system’s vast knowledge by envisioning possible and very desirable output. In what follows, we discuss specific cognitive gaps in envisioning LLM interactions:
3.2.1. Capability Gap:
Recall from our earlier definition that an intention includes specifications about how to execute task processes. The capability gap concerns the users’ inability to formulate “how to” procedures to implement their intentions. Defining when, where, and how one wants to take action on intentions requires added cognitive effort (Patalano and Seifert, 1997; Seifert and Patalano, 2001), but has been shown to enhance the rate of successful execution (Gollwitzer and Oettingen, 2020). While LLMs, with their extensive training data, can theoretically understand and generate a wide variety of bespoke task content, their very strength can also be a source of complexity for users. Take, for instance, entering a prompt for summarizing a 2-page text in 1 page. For a human writer, this task involves a myriad of choices about what to prioritize, which nuances to retain, and which details to omit. How can a user identify the right specification of intention for the LLM?
For most generation tasks, a cognitive process is not already well-formed in memory but can be made more explicit (or newly generated) through planning. As described in the section on cognitive models, generative tasks are not determinate and vary with each instance of generation and each intention (every story is written with a different specific process). On the other hand, when tasks are well practiced through experience with execution, people may have “scripts” or well-developed routines to complete a generative task  (Schank and Abelson, 2013). However, some tasks may be too well learned to summarize as instructions; for example, asking someone how to make a peanut butter and jelly sandwich without actually making it [^1] The EXACT INSTRUCTIONS CHALLENGE video: https://youtu.be/Ct-lOOUqmyY demands generating actions without feedback from execution. “Doing” offers reminders of where you take added steps to avoid errors. Without this step (since LLMs are doing the generation), users are forced to take a trial-and-error approach with differing abstractions and variations of their intentions, and then iterate based on evaluating the LLM’s output. But unlike deterministic systems where a given input always produces the same output, an LLM can produce countless variations from the same input, and there is, by definition, no one correct answer in generative tasks. This indeterminacy in LLM responses, coupled with a lack of transparency in how inputs lead to LLM outputs, can hinder users’ attempts to formulate prompts.
Of course, human-LLM interactions are designed to be conversational and iterative, allowing users to evaluate output, revise prompts, and further refine intentions until satisfied. However, there are costs associated with iteration from user experience, social exchange expectations (Cropanzano and Mitchell, 2005), and generative process perspectives. In envisioning intentions, users must find a balance between “batching” intentions in prompts and development through iteration. Intention setting in advance allows users to carefully consider their goals without potential change or interference from the LLM. For example, if writing about personal experience or emotions, iterations may draw the eventual outcome further away from the user’s internal perspective. Repeated iteration over many small changes to prompts may violate a user’s conversational expectations for “chunking” of information in turn-taking. Most importantly, iteration poses the danger of fixation on the presented output. Humans tend to become fixated on their own or others’ initial proposed solutions  (Maier, 1931; Jansson and Smith, 1991; Leahy et al., 2020) and struggle with overcoming attachment to early ideas  (Cross, 2001), limiting options for solutions through fixation on a particular type of idea or concept (Purcell and Gero, 1996), resulting in a lack of exploration of alternatives. If a first output is viewed, the user may “anchor” on it as a solution and focus on local refinements to local minima, thereby missing opportunities in more distant solution spaces. Those with less developed intentions may be more likely to show a stronger fixation on an initial LLM output due to a lack of knowledge about desirable outcomes. Experienced users may learn when greater effort to formulate intentions within initial prompts is worthwhile in arriving at better solutions or quicker system interactions. For example, while it is possible to iterate based on output evaluation, the time to read and reformulate prompts has costs. If the iterative steps can be combined in fewer steps, taking more care in formulating intentions may be cost-effective.
3.2.2. Instruction Gap:
The instruction gap refers to the user’s challenges in clearly and effectively expressing their intentions in the interface as text prompts, leading to a potential mismatch between the user’s intentions and what the LLM perceives and then produces. One benefit to the natural language interface for LLMs for end-users is a wide-open input space for specifying intentions. In writing the toast, the user’s intentions as described in the text prompt can vary across language expressions; for example, the role assumed (e.g., “act as an event emcee,” “act as a best friend”), desirable qualities of output (e.g., “make it funny,” “make it heartfelt”), and related text content as a starting point for generation (e.g., “golf,” “Vegas”). Unfortunately, the LLMs’ algorithms for learning from text corpora create a dependency on language precision. While human language use tolerates a wide range of expressions communicating a similar meaning, even slight changes in words can lead LLMs to produce significantly different outputs. It is possible that substantial experience with LLM use will improve a user’s understanding of the precise mapping between prompt contents, at least in a domain area. For instance, more expert users may have an understanding of transformer architectures, and experienced users may know how specific keywords in the prompt (e.g., “punchy”) influence the attention mechanism. “Instructing” is a much more deliberate task with individual variation, and there are few examples of product instructions done well despite their ubiquity (but see  (Agrawala et al., 2003)). Users challenged in performing improvisational linguistic dexterity may not obtain satisfactory outcomes solely through trial and error iteration of prompts. Further, linguistic sensitivity makes it difficult for users to script their intentions for the LLM upfront by anticipating potential system interpretations. Therefore, users are likely to need a multi-turn interaction with an LLM session to determine alternative specifications for their intentions leading to better outcomes.
@@ 3.2.3. Intentionality Gap:
In PATH-1 interaction with the LLM, the user states their goal directly and forgoes any cognitive task processes anticipated in doing the tasks themselves (i.e., planning what to write, executing it, and evaluating it). Instead, they start with evaluation of the LLM-generated text. Does it meet their goal? This evaluation is cognitively demanding because no “bridging” steps between goal and output are provided. By skipping directly from goal to evaluating output, users may lack a cohesive understanding of the task context and required content, hindering their ability to make comprehensive judgments about the adequacy of the output. Drawing from philosophical terminology, users encounter an intentionality gap while assessing texts generated by LLMs. In philosophy, intentionality refers to the capacity of mental states to be about, directed at, or represent something in the world outside of themselves (Searle, 1983). In the LLM context, users may not have a nuanced “aboutness” of the specific intentions or contexts they are considering and, therefore, may not have a clear sense of what the content should be “about” from their own perspective. Consequently, without internal cognitive benchmarks of the generation process for comparison, they may become over-reliant on the LLM and avoid evaluation against thier self-generated goals. Further, engaging in subsequent planning to identify and rectify any deviations from a desired outcome can be challenging. In other words, users producing only a goal statement will have difficulty answering questions such as, “Can the LLM do better?” and “How can the LLM do better?”
@@
Note that this intentionality gap is different from the gulf of evaluation in Norman’s model (Norman, 1986). With traditional HCI tasks, deterministic outcomes are easier to perceive and interpret, whether completed successfully or not. Feedback is often immediate, as with a direct manipulation user interface in which “…users can immediately see if their actions are furthering their goals (Shneiderman, 1982).” The evaluation gulf stems from ambiguous system feedback at the interface level, which may make it challenging to perceive and interpret the produced outcome. Addressing the gulf of evaluation involves improving system feedback, representation, and visibility. In contrast, the intentionality gap lies at the cognitive level due to the user’s failure to create clear intentions. Bridging this intentionality gap with LLMs may require interfaces to scaffold users in creating clearer task process prompts and developing their contextualized understanding of the qualities desired in the generated text.
The idea that under-specified intentions lead to challenges in evaluation surfaces in Karl Duncker’s (Duncker, 1945) work on the process of finding a solution through a continual restructuring of a problem over time to develop the “essential” properties of a desired solution. Problems intentionally left open-ended, as in many creative tasks, are termed “ill-defined” (Simon and Newell, 1971) in that incomplete information is provided about the problem. For example, design intentions require a great deal of construction by the designer  (Restrepo and Christiaans, 2004) to identify valuable qualities of potential solutions. Identifying the nature of a problem is key to solving problems, creative work, and design thinking  (Hargadon and Bechky, 2006; Dorst, 2011; Ecker, 1963). Exploring problems provides perspectives on values important in solutions (Dorst and Cross, 2001; Schon, 1983). The process of developing intentions helps one learn to “see” the desirable attributes of non-existent outcomes (Dorst, 2011) and specify important solution qualities. The power of exploring intention is illustrated by an early study of fine artists asked to create still-life drawings: those showing “discovery-oriented” behavior (rearranging objects, changing perspectives, touching objects) produced works with greater originality and higher quality (as judged by experts), and even experienced greater professional recognition and income years later. Those who showed more consideration of the qualities of their intended piece produced better outcomes than those focused on execution (Csikszentmihalyi and Getzels, 1971, 1988).